OpenStudy (anonymous):

Take the derivative of \[(d/dx )\int\limits_{a}^{\cos(x^(2)} \arcsin(\ln(t + 1))dt\] would the answer be arcsin(ln(-2xcos(x^(2))(sin(x^(2)) + 1)) ? or arcsin(ln(cos(x^(2)) + 1)*(-2xcos(x^(2))(sin(x^(2)) + 1))

OpenStudy (anonymous):

b = cos(x^(2)) sorry about the mistake in latex

OpenStudy (anonymous):

sorry is this the answer? arcsin(ln(cos(x^(2)) + 1))*(-2xcos(x^(2))(sin(x^(2)) + 1))

OpenStudy (anonymous):

Im assuming that it would be b = cos(x^(2)) thus arcsin(ln(b+1)*b')

myininaya (myininaya):

\[\text{ Let } f(t)=\arcsin(\ln(t+1))\] Let F be the antiderivative of f =>(F)'=f So we have \[\frac{d}{dx}[\int\limits_{a}^{\cos(\frac{x}{2})}f(t) dt ]=\frac{d}{dx}[ F(t)|_a^{\cos(\frac{x}{2})}]\] \[\frac{d}{dx}[F(\cos(\frac{x}{2}))-F(a)]=\frac{d}{dx}F(\cos(\frac{x}{2}))-\frac{d}{dx}[F(a)]\] \[\text{ apply chain rule } (\cos(\frac{x}{2}))'f(\cos(\frac{x}{2}))-0=(\cos(\frac{x}{2}))'f(\cos(\frac{x}{2})\]

myininaya (myininaya):

\[\frac{1}{2}(-\sin(\frac{x}{2})) \cdot \arcsin(\ln(\cos(\frac{x}{2})+1))\]

myininaya (myininaya):

Well looking at what @cuddlepony said I don't know if I have the right limits

OpenStudy (anonymous):

you are suppose to ignore a

OpenStudy (anonymous):

you are simply canceling the integral

myininaya (myininaya):

lol oh you are the one asking the question i'm so dumb

OpenStudy (anonymous):

pff no you are not we all make mistakes

OpenStudy (anonymous):

whats right? is arcsin(ln(b+1)*b') the correct answer?

myininaya (myininaya):

Well the problem I put down is a good example for yours

myininaya (myininaya):

He is talking about the answer to the problem I thought we had

OpenStudy (experimentx):

i mean the process.

OpenStudy (anonymous):

the process is you do a substitution of b = cos(x^(2)) then take the derivative with respect to b it cancels the integral then you just apply chain rule to get arcsin(ln(b+1)*b') I'm confused as to where b' goes but i guess outside of the ln(b+1) seems correct